OpenStudy (anonymous):

Establish this identity: tan(θ)cot(θ)-sin^2(θ) = cos^2(θ)

jimthompson5910 (jim_thompson5910):

hint: cot(x) = 1/tan(x) and sin^2(x) + cos^2(x) = 1

OpenStudy (luigi0210):

Easy, what's tan and cot in terms of sin/cos?

OpenStudy (anonymous):

did you eat magic beans?

OpenStudy (luigi0210):

Me and Jim have different methods, but same result

OpenStudy (anonymous):

Gah. I'll come back to this tomorrow, I've been doing homework for way to long. I'll look at it again tomorrow. Good night. Lol.

OpenStudy (anonymous):

it would be easier if you study "The Function Hexagon". It will make trigonometry simple. :) tangent and cotangent are inverse functions right? so therefore it is equal to 1 when multiplied to each other. now you have: 1-sin^2(theta) = cos^2(theta) look back again the formula, sin^2(theta)+cos^2(theta) = 1 if we transpose "sin^2(theta)" to the other side, it will be: cos^2(theta) = 1-sin^2(theta) right? now you proved it. cos^2(theta) = cos^2(theta)

OpenStudy (anonymous):

Oh okay I see what you did. and the tan was like adding fractions . . . cot and the tan was like adding fractions . . . Brain fart. Thank you so much for your help. :)